A chemist measures the enthalpy change

2Na

(s)

Cl2

(g)

2NaCl

(s)

/kJmol

.

reaction

12Na

(s)

14Cl2

(g)

12NaCl

(s)

2NaCl

(s)

2Na

(s)

Cl2

(g)

4NaCl

(s)

4Na

(s)

2Cl2

(g)
<div id="ansed_tree_ansedimage_27" class="ansed_tree" overflow:visible;"="" unselectable="on" aria-hidden="true" style="position: absolute; left: 0px; top: 0px; z-index: 3; width: 32px; height: 24px;"> kJ
Answer #1

Entropy change = no . Of molecules of product - no. Of reactant molecules .

1. 2NaCl - (2Na + 1 Cl2) = -1 .

2. 12- 26 = -14

3. 2-3 = -1

4 . 4- 6 = -2 .